OpenStudy (anonymous):

4. One of the offices at the company was being repainted. Riley’s father challenged him to figure out how many quarts of paint would be needed. The dimensions of the office are 16 ft by 12 ft by 10 ft. The office floor, ceiling, and door would not be painted and there are no windows in this office. The door opening measures 3 ft by 8 ft. One quart of paint covers 100 ft2. How many quarts of paint would be needed to paint the room, if only full quarts can be purchased? Show your work.

OpenStudy (danjs):

First , figure out the surface area of the 4 walls of the office. Then, subtract out the door area.

OpenStudy (danjs):



OpenStudy (danjs):

two walls are 12 by 10, and two walls are 16 by 10

OpenStudy (danjs):

The total area of the 4 walls = (2 small wall areas) + (2 larger wall areas) = 2*(12*10) + 2*(16*10)

OpenStudy (danjs):

right?

OpenStudy (danjs):

where you go

OpenStudy (anonymous):

560

OpenStudy (danjs):

k, that is the total area of the 4 walls, need to take out the area of the door that will not be painted... what is the area of the door? A= length x width

OpenStudy (anonymous):

24

OpenStudy (danjs):

yep, Area to be painted = Area of 4 walls - Area of the door

OpenStudy (anonymous):

536

OpenStudy (danjs):

ok, each quart of paint covers 100 ft^2 of area

OpenStudy (danjs):

\[\frac{ 536 ~ft^2 }{1 }*\frac{ 1~quart~paint }{ 100~ft^2 } = ? ~~Quarts Paint\]

OpenStudy (danjs):

Then round to nearest whole number of quarts, since the problem says, "only full quarts can be purchased"

OpenStudy (danjs):

good?

OpenStudy (anonymous):

no

OpenStudy (danjs):

536 divided by 100 = 5.36 quarts of paint needed The paint only comes in full quarts, so you need to buy 6 quarts

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank u